OpenStudy (anonymous):

Calculate the average rate of change for the given function, from x = −3 to x = 7.

OpenStudy (anonymous):



OpenStudy (anonymous):

one fifth negative one fifth −5 5

OpenStudy (anonymous):

(0-50)/(7- (-3)) = -50/10 = -5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome.